One number is 8 less than another. If 5 times the smaller number minus 2 times the larger number is 14, find the two numbers.
Answer by checkley75(3666)

x-y=8, x=y+8
5y-2x=14
5y-2(y+8)=14
5y-2y-16=14
3y=14+16
3y=30
y=30/3
y=10 answer.
x-10=8
x=8+10
x=18 answer.
proof
5*10-2*18=14
50-36=14
14=14